HETAV PATEL

Grade 11,

(a+b)^2 pol,mnhjknbvdrthbvcxsertybvcdftyujnbvcdfrtyujnbvcdftyjnbvfrtyujbvftyunbvftyujnbg

(a+b)^2 pol,mnhjknbvdrthbvcxsertybvcdftyujnbvcdfrtyujnbvcdftyjnbvfrtyujbvftyunbvftyujnbg

Grade:12th pass

3 Answers

vijay kale
6 years ago
This is a very popular algebraic identity (a+b )^2 =a^2+b^2+2ab. This can be easily proved with distributive proverty.
Om
5 years ago
Dear student
Please check whether you wrote the correct question or not
I hope you explain your question in an understanding way
 
With regards
Your Fellow tudent
Sayantan Garai
5 years ago
(a+b)²=a²+2ab+b²